# Chaotic discretization theorems for forced linear and nonlinear coupled oscillators

### v2

@@ Abstract
We prove the holding of chaos in the sense of Li-Yorke for a family of four-dimensional discrete dynamical systems that are naturally associated to ODEs systems describing coupled oscillators subject to an external non-conservative force, also giving an example of a discrete map that is Li-Yorke chaotic but not topologically transitive. Analytical results are generalized to a modular definition of the problem and to a system of nonlinear oscillators described by polynomial potentials in one coordinate. We perform numerical simulations looking for a strange attractor of the system; furthermore, we perform a bifurcation analysis of the system presenting 1D and 2D bifurcation diagrams, together with spectra of Lyapunov exponents and basins of attraction.
@@ 1 Introduction
It is worth to mention some recent results. In Bernardes Jr et al. [2015] Li-Yorke chaos is analytically studied for linear operators on Banach and Fréchet spaces, while Li and Wang [2025] deeply focus on a stronger definition of dense uniform Li-Yorke chaos. In Zhang et al. [2023] it is shown how Li-Yorke chaos depends on the topology defining the dynamical system. The topic has been explored also for PDEs Zhang et al. [2025]. See Garić-Demirović et al. [2024], Wei and Zhou [2019] for explicit derivations of Li-Yorke chaos for specific models. An interesting emerging field concerns chaotification techniques for the so-called “anti-control of chaos”; see Moysis et al. [2025] for an exhaustive review.
@@
The paper is organized as follows. In Section 2 we recall some basic definitions together with the Li-Yorke theorem and the Marotto-Li-Chen theorem Marotto [1978], Li and Chen [2003a]; then, we define the discrete map subject of this work together with a simple notion of discretization. In Section 3 we prove that the system is not topologically transitive for a symmetric choice on the parameters. In Section 4 we prove the emergence of Li-Yorke chaos for proper choices on the parameters and generalize results for a modular definition of the system, leading to three chaotic discretization theorems for forced linear and nonlinear coupled oscillators. In Section 4.1 we comment a possible connection between Li-Yorke chaos and the KAM theory Kolmogorov [1979], Arnold [1963], Moser [1966] for quasi-integrable systems. In Section 5 we present some numerical results. In Section 5.1 we perform qualitative studies of the orbit of the system, by comparing regular orbits for a symmetric choice on the parameters and the rising of a confined strange attractor in the general case. In Section 5.2 we present 1D and 2D bifurcation diagrams of the system and confirm the rising of period doubling bifurcations through a bifurcation analysis using the MATLAB tool MatContM (https://matcont.sourceforge.io/). Furthermore, we present spectra of Lyapunov exponents, that confirm its sensitive dependence on initial data for several range of the parameters, together with basins of attraction of the system. Finally, in Section 6 we state our conclusions and propose further developments of this work.
@@ 2 Mathematical preliminaries
We stress the fundamental consequence of the Li-Yorke theorem: any one-dimensional map defined on compact sets that has a periodic point of period 3 will show itself as chaotic in the sense of Li-Yorke, i.e. it will admit a uncountable set of points that are simultaneously close and distant in the sense of Definition 1.
@@
Notice that the discretization notion we define is not a standard discretization algorithm for ODEs systems, such as explicit or implicit Euler discretization (see, e.g., Kazakov et al. [2024]), and in general it is not appropriate for the numerical resolution of the associate ODE system. For instance, we can easily prove that for Hamiltonian systems it does not preserve in general the symplectic structure of the phase space, as shown in Proposition 1.
@@
where J is the Jacobian of the system (2). A simple calculation is sufficient to see that the symplecticity condition holds only for the specific class of Hamiltonian that satisfy
@@
Nevertheless, our aim is not to obtain a discrete approximation of the ODEs system, but rather to define an alternative discrete version associated to the continuous system and to explore connections between regular and chaotic behavior of both systems. The discrete system can share some properties with the continuous one (e.g. fixed points); on the other hand, it can exhibit a richer dynamics. Furthermore, the discretization we define can be thought as a novel chaotification procedure that associates a Li-Yorke chaotic discrete system to an integrable continuous one, under suitable constraints on the parameters.
@@ 4 Chaotic discretization theorems
where $\varepsilon_{1,2}>0$, $\alpha\neq 0$, $\gamma\in\mathbb{R}$
@@ 4.1 Qualitative remarks on scrambled sets and perturbations
In answering this question, a curious connection with the KAM theorem Kolmogorov [1979], Arnold [1963], Moser [1966] can be conjectured. From KAM theory, we know that a quasi-integrable system exhibits invariant tori in the phase space that are not destroyed by the presence of a perturbation; when acting on the system, the perturbation destroys some tori, while other ones survive but are deformed, according to specific non-resonance conditions. In the context of chaos theory, it might be argued that a perturbation acts on the scrambled set of a Li-Yorke chaotic map in a similar way, in the sense that they could exist specific sufficient conditions such that the scrambled set of a Li-Yorke map survives under a small perturbation. If this is true, it would be possible to deduce chaos for a discrete system in general through perturbative arguments.
@@ 5.1 Looking for a strange attractor
Figure 1 refers to the symmetric case $\tilde{\alpha}=-\alpha$; we can notice the very sensitive dependence of the system on initial data.
@@ 6 Conclusions
In this work we analyzed a family of four-dimensional discrete dynamical systems defined through polynomial and periodic functions of one coordinate, proving that the map is chaotic in the sense of Li-Yorke under suitable requests on the parameters. Relevance of Theorem 3 - 5 lies on the connection that it establishes between chaotic behavior of these discrete maps and respective continuous systems, when discretized in a proper way. Discrete maps satisfying hypothesis of Theorem 3 - 5 can be seen as discrete versions of several continuous dynamical systems, that can be integrable, quasi-integrable or chaotic; regardless of the regularity of the motion exhibited by these systems, their discrete version, as considered in this work, is always Li-Yorke chaotic. This fact is open to some interpretations. In particular, it shows how the necessary conditions for the emerging of chaos strongly depend on the choice between a discrete or continuous evolution step, at least within the framework considered in this work; on the other hand, it is a further confirmation of the generality of the definition given by Li-Yorke, that is sufficiently weak to be satisfied by a large class of dynamical systems.
@@
The rising of a strange attractor for the system leads to some questions. Firstly, we might numerically compute the Hausdorff dimension Hausdorff [1919] of the attractor, in order to have a confirmation of its fractal dimension. Secondly, it is not immediate to provide an explanation for certain dynamical properties, such as the rising of the attractor only on certain projection planes or the peculiar regular orbits rising in the symmetric case. Finally, a deeper bifurcation analysis of the system may be performed, in order to find further bifurcation points and provide a complete scenario for stability and transition to chaos for the system.
@@
During the preparation of this work the author used Google Gemini (Large Language Model) in order to write the Python code for reproducing the 2D bifurcation diagrams and the Lyapunov exponents spectra (Figure 16 - 21). After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.

### v3

@@ Abstract
We prove the holding of chaos in the sense of Li-Yorke for a family of four-dimensional discrete dynamical systems that are naturally associated to ODE systems describing coupled oscillators subject to an external non-conservative force, also giving an example of a discrete map that is Li-Yorke chaotic but not topologically transitive. Analytical results are generalized to a modular definition of the problem and to a system of nonlinear oscillators described by polynomial potentials in one coordinate. We perform numerical simulations looking for a strange attractor of the system; furthermore, we perform a bifurcation analysis of the system presenting 1D and 2D bifurcation diagrams, together with spectra of Lyapunov exponents and basins of attraction.
@@ 1 Introduction
It is worth to mention some recent results. In Bernardes Jr et al. [2015] Li-Yorke chaos is analytically studied for linear operators on Banach and Fréchet spaces, while Li and Wang [2025] deeply focuses on a stronger definition of dense uniform Li-Yorke chaos. In Zhang et al. [2023] it is shown how Li-Yorke chaos depends on the topology defining the dynamical system. The topic has been explored also for PDEs Zhang et al. [2025]. See Garić-Demirović et al. [2024], Wei and Zhou [2019] for explicit derivations of Li-Yorke chaos for specific models. An interesting emerging field concerns chaotification techniques for the so-called “anti-control of chaos”; see Moysis et al. [2025] for an exhaustive review.
@@
The paper is organized as follows. In Section 2 we recall some basic definitions together with Li-Yorke theorem and Marotto-Li-Chen theorem Marotto [1978], Li and Chen [2003a]; then, we define the discrete map subject of this work together with a simple notion of discretization. In Section 3 we prove that the system is not topologically transitive for a symmetric choice on the parameters. In Section 4 we prove the emergence of Li-Yorke chaos for proper choices on the parameters and generalize results for a modular definition of the system, leading to three chaotic discretization theorems for forced linear and nonlinear coupled oscillators. In Section 4.1 we comment a possible connection between Li-Yorke chaos and KAM theory Kolmogorov [1979], Arnold [1963], Moser [1966] for quasi-integrable systems. In Section 5 we present some numerical results. In Section 5.1 we perform qualitative studies of the orbit of the system, by comparing regular orbits for a symmetric choice on the parameters and the rising of a confined strange attractor in the general case. In Section 5.2 we present 1D and 2D bifurcation diagrams of the system and confirm the rising of period doubling bifurcations through a bifurcation analysis using the MATLAB tool MatContM (https://matcont.sourceforge.io/). Furthermore, we present spectra of Lyapunov exponents, that confirm sensitive dependence on initial data for several range of the parameters, together with basins of attraction of the system. Finally, in Section 6 we state our conclusions and propose further developments of this work.
@@ 2 Mathematical preliminaries
We stress the fundamental consequence of the Li-Yorke theorem: any one-dimensional map defined on intervals that has a periodic point of period 3 will show itself as chaotic in the sense of Li-Yorke, i.e. it will admit a uncountable set of points that are simultaneously close and distant in the sense of Definition 1.
@@
Notice that the discretization notion we define is not a standard discretization algorithm for ODE systems, such as explicit or implicit Euler discretization (see, e.g., Kazakov et al. [2024]), and in general it is not appropriate for the numerical resolution of the associate ODE system. For instance, we can easily prove that for Hamiltonian systems it does not preserve in general the symplectic structure of the phase space, as shown in Proposition 1.
@@
where J is the Jacobian of the system (2). A simple calculation is sufficient to see that the symplecticity condition holds only for the specific class of Hamiltonians that satisfy
@@
Nevertheless, our aim is not to obtain a discrete approximation of the ODE system, but rather to define an alternative discrete version associated to the continuous system and to explore connections between regular and chaotic behavior of both systems. The discrete system can share some properties with the continuous one (e.g. fixed points); on the other hand, it can exhibit a richer dynamics. Furthermore, the discretization we define can be thought as a novel chaotification procedure that associates a Li-Yorke chaotic discrete system to an integrable continuous one, under suitable constraints on the parameters.
@@ 4 Chaotic discretization theorems
where $\varepsilon_{1,2}>0$, $\alpha\neq 0$, $\gamma\in\mathbb{R}$,
@@ 4.1 Qualitative remarks on scrambled sets and perturbations
In answering this question, a curious connection with the KAM theorem Kolmogorov [1979], Arnold [1963], Moser [1966] can be conjectured. From KAM theory, we know that a quasi-integrable system exhibits invariant tori in the phase space that are not destroyed by the presence of a perturbation; when acting on the system, the perturbation destroys some tori, while others survive but are deformed, according to specific non-resonance conditions. In the context of chaos theory, it might be argued that a perturbation acts on the scrambled set of a Li-Yorke chaotic map in a similar way, in the sense that they could exist specific sufficient conditions such that the scrambled set of a Li-Yorke map survives under a small perturbation. If this is true, it would be possible to deduce chaos for a discrete system in general through perturbative arguments.
@@ 5.1 Looking for a strange attractor
Figure 1 refers to the symmetric case $\tilde{\alpha}=-\alpha$, where we can notice the very sensitive dependence of the system on initial data.
@@ 6 Conclusions
In this work we analyzed a family of four-dimensional discrete dynamical systems defined through polynomial and periodic functions of one coordinate, proving that the map is chaotic in the sense of Li-Yorke under suitable requests on the parameters. Relevance of Theorem 3 - 5 lies on the connection that they establish between chaotic behavior of these discrete maps and respective continuous systems, when discretized in a proper way. Discrete maps satisfying hypothesis of Theorem 3 - 5 can be seen as discrete versions of several continuous dynamical systems, that can be integrable, quasi-integrable or chaotic; regardless of the regularity of the motion exhibited by these systems, their discrete version, as considered in this work, is always Li-Yorke chaotic. This fact is open to some interpretations. In particular, it shows how the necessary conditions for the emerging of chaos strongly depend on the choice between a discrete or continuous evolution step, at least within the framework considered in this work; on the other hand, it is a further confirmation of the generality of the definition given by Li-Yorke, that is sufficiently weak to be satisfied by a large class of dynamical systems.
@@
The rising of a strange attractor for the system leads to some questions. Firstly, we might numerically compute the Hausdorff dimension Hausdorff [1919] of the attractor, in order to have a confirmation of its fractal dimension. Secondly, it is not immediate to provide an explanation for certain dynamical properties, such as the rising of the attractor only on certain projection planes or the peculiar regular orbits rising in the symmetric case. Finally, the bifurcation analysis performed in this work may be extended to larger sets of parameters, in order to provide a complete scenario for stability and transition to chaos for the system.
@@
During the preparation of this work the authors used Google Gemini (Large Language Model) in order to write the Python code for reproducing the 2D bifurcation diagrams and the Lyapunov exponents spectra (Figure 16 - 21). After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.